OpenStudy (anonymous):

Solve M = 2x + 3y for y. (1 point)

OpenStudy (saifoo.khan):

(M-2x)/3 = y

OpenStudy (anonymous):

y=(M-2x)/3 y=(M/3)-(2x/3)

OpenStudy (anonymous):

(M-2x)/3=y